Circular No. 6102

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

ROSAT SOURCE CATALOGUE H.-U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik

(MPE), Garching, writes: “The ROSAT Consortium (MPE, Garching; Goddard Space Flight Center; Smithsonian Astrophysical Observatory; and the WFC-Consortium in the U.K.) announces the public release on Oct. 26 of The First Rosat Source Catalogue of Pointed Observations with the PSPC. This catalogue contains 50408 sources from 2876 pointed observations. For each source, the following properties are provided: the observation number, the ROSAT name, the position in equatorial and in galactic coordinates, four positional errors (intrinsic, systematic, boresight, and total), the source countrate and its error, the background countrate, exposure time, hardness-ratios HR1 and HR2 and their errors, extent and likelihood of extent, likelihoods of the map-detect algorithm and of the maximum-likelihood detection algorithm, etc. The present version of the catalogue comprises observations that have been performed until the end of May 1993 and which are in the public ROSAT Data Archive. Confused sources and 124 observations with overly crowded regions or regions of highsurface-brightness diffuse emission have been removed. The resulting 50408 sources have a likelihood of at least 10, corresponding to a rate of accidental detections of ∼ 1 percent. The new catalogue can be accessed via the interactive account (telnet) xray.mpe-garching.mpg.de with username xray (no password) by typing browse rosatsrc. Database ‘rosatseq’ (browse rosatseq) contains information on the individual observation fields. Questions/comments may be e-mailed to rosat svc@mpe-garching.mpg.de.”

RE J1255+266 W. Wenzel, Sonneberg Observatory, communicates: “No trace of RE

J1255+266 (cf. IAUC 6085) appears to a threshold of m_{pg} = 11.5–12 on 16 sky-patrol plates taken by B. Fuhrmann between Jan. 17 and June 7.”

HD 5980 F. M. Bateson, Variable Star Section, Royal Astronomical Society of

New Zealand, reports (via A. Gilmore): “HD 5980 (cf. IAUC 6099) has been monitored by A. F. Jones, Nelson, since 1958. He first detected an increase in brightness of 1 mag on 1994 July 18.7 UT (VSS Circular M94/7, RASNZ).”

1994 November 9 Daniel W. E. Green